OpenStudy (anonymous):

\[2+3i \div-1-4i\] simplified

OpenStudy (anonymous):

\[\frac{2+3i}{-1-4i}\] gimmick is to multiply top and bottom by conjugate of the denominator, which is \[-1+4i\]

OpenStudy (anonymous):

reason being that \[(a+bi)(a-i)=a^2+b^2\] a nice real number with no i in it

OpenStudy (anonymous):

so write \[\frac{2+3i}{-1-4i}\times \frac{-1+4i}{-1+4i}\]

OpenStudy (anonymous):

you denominator will be \[1^2+4^2=17\] and your numerator will be whatever you get when you multiply

OpenStudy (anonymous):

okay thanks!

OpenStudy (anonymous):

wait would a negative or positive 14 be on the top?